Home‎ > ‎Crew 169 Team Peru IV‎ > ‎

0509-DailySummaryReport

DAILY SUMMARY REPORT

SOL: (8)

Person filling out Report: XO

Summary Title: Phase IIIa.

Mission Status: Culture on Rover recovery

Sol Activity Summary: Today I’m realizing a sampling on principal parts of Rover recovery before it EVA.

Look Ahead Plan: The process was a sampling of wheel and the principal cube robot. These samples were isolate in the three Petrifilm medium.

Anomalies in work: No problems.

Weather: Phase IIIa, sampling and isolate.

Crew Physical Status: Adequate.

EVA: Initial (1).

Reports to be file: Commander Report, XO Report, Engineering Report, HSO report.

Support Requested: distilled water, stereoscope and cotton tipped.